OpenStudy (anonymous):

\[\sqrt[4]{y-4?}=3\] solve

OpenStudy (anonymous):

85

OpenStudy (anonymous):

y-4=3^4 y=85

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

y=85